OpenStudy (anonymous):

Check my intervals of increase and decrease for function f(x)=2radical(x)-x. I got Dec:(-infinty,-1), (-1,1) Inc: (1,2),(2,infinty)

OpenStudy (shubhamsrg):

what is radical(x) ??

OpenStudy (anonymous):

Shouldn't the inc be (2,4) not (2,infinity)

OpenStudy (anonymous):

\[f(x)=2\sqrt{x}-x \] is the equation

OpenStudy (shubhamsrg):

whats your f'(x) ?

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{x} }-1\]

OpenStudy (shubhamsrg):

correct equating that to 0, you'll get x=1 as the only critical point, right ?

OpenStudy (anonymous):

Yup

OpenStudy (shubhamsrg):

now, whats your f''(x)

OpenStudy (anonymous):

\[\frac{ -1 }{2x ^{3/2} }\]

OpenStudy (anonymous):

so tiny, its ^3/2

OpenStudy (shubhamsrg):

yeps, f''(1) <0 , that means x=1 should be a point of maxima, right ?

OpenStudy (anonymous):

yesss. thats what i got

OpenStudy (shubhamsrg):

Also, note that your function is continuous everywhere .

OpenStudy (shubhamsrg):

Hence, it must be inc in the interval (-inf,1) and dec in (1,inf) at x=1, its neither inc not dec.

OpenStudy (shubhamsrg):

Any problems ? hope I didn;t make any mistake ?

OpenStudy (shubhamsrg):

oh am so sorry, I overlooked one thing

OpenStudy (anonymous):

Im a little confused about finding the intervals, dont you pick a number in between 1 and something bigger like 2. so youd plug 1/2 into the first derivative to find it?

OpenStudy (shubhamsrg):

x<0 is not a part of the domain.

OpenStudy (anonymous):

To find if its increasing or dec. sorry

OpenStudy (shubhamsrg):

see, f'(x) represents slope of the function. f'(x) = 0 means we are finding that point where slope is 0, i.e. tangent is parallel to x axis. we found x=1 is the only point there. checking for f''(x), we find f''(1) < 0 which means rate of change of slope at x=1 is negative, or in other words, x=1 should be a point of maxima the rough graph of the function must be something like this :  this signifies left of 1, graph should be increasing, to its right, it must be decreasing.

OpenStudy (shubhamsrg):

Am not following exactly what your query is ? the answer should be increasing in [0,1) , and decreasing in (1, inf) p

OpenStudy (anonymous):

No that makes sense! I can look at the graph and see that. But i cant find it using the critical number if that makes sense.

OpenStudy (shubhamsrg):

you are trying to say, that you find f(1) , then you take any other point, say 2, and find f(2) , if f(1) > f(2) , function is deceasing to the right of 1 , same wahy f(1/2) < f(1) , so func is inc to the left of 1 ?? still not unable to understand your query ?